OpenStudy (tennistar):

The area of a rectangle is the product of width and length. If a rectangle has length 15mthe square root of 13 times m. and width the square root of 9 times m., which of the following is the area of the rectangle?

OpenStudy (tennistar):

I got now what?

OpenStudy (anonymous):

Is this the problem? Find the area if: length = \(15m\sqrt{13}m\) width = \(\sqrt{9}m\)

OpenStudy (tennistar):

yes

OpenStudy (anonymous):

Can you possibly draw the problem as it is or if it's on the computer, screenshot it?

OpenStudy (anonymous):

\[A = 15m\sqrt{13 \times 3^{2} \times m^{2}}\]Can you solve it now? Just take the squared parts outside of the radical.

OpenStudy (tennistar):

so a = 15m * 3^2 8m^2 sqrt13

OpenStudy (anonymous):

Well, what's the \(\sqrt{9}\)?

OpenStudy (tennistar):

please just give me an answer, I am on a time limit

OpenStudy (anonymous):

Sorry, but it's against CoC. Actually, we would've been done a minute ago if you just answered that, but here's the next step. \[15(3)(m)(m)\sqrt{13}\]